OpenStudy (fools101):

Logarithmc Equations log3 x - log3 4 = 2

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

log 3 (x/4) = 2 3^2 = x/4 9=x/4 x = 36

OpenStudy (anonymous):

i assume u meant base 3

OpenStudy (fools101):

yes! and thank u Mathwiz! :)